OpenStudy (anonymous):

0.4x + 8 is less than or equal to 1.2x-7 x (inequality sign) = ?

OpenStudy (anonymous):

0.4x + 8 \[\le\] 1.2x-7 0.4x-1.2x\[\le\]-7-8 -0.8x\[\le\]-15 x\[\le\]15/0.8 x\[\le\]18.75